i don't understand solving these types of equations...
x-5=-5 What does x equal?
0=y-2 What does y equal?
2y to the power of 2 = 0
8=y + 3y
2x to the power of 2= 0
Can anyone help me??

Answer by emmychan(11)

These types of equations are confusing for people because since when do letters belong in math? Think of it like this...
Every letter is only a symbol. It's like when your teacher puts (?) and you have to solve. Try putting a (?) and solve.
x - 5 = -5
(?) - 5 = -5 (Now add 5 to both sides)
+ 5 + 5
(?) = 0 (put x back)
x = 0

0 = y - 2
0 = (?) - 2 (Add 2 to both sides)
+ 2 + 2
2 = (?) (Put y back)
2 = y